OpenStudy (anonymous):

A bank teller has 35 bills in $5,$10, and $20 denominations. The total value from these bills is $370. The total number of $10 bills and $20 bills is 7 more than the number of $5. How many bills of each denomination does the teller have?

OpenStudy (anonymous):

The first thing that I'm doing here, which I'm not sure if its on the right track, but: I looked for two numbers with a difference of 7 that when added up will give you 35, since that is the total number of bills and the problem states that "the total number of 10 bills and 20 bills is 7 more than 5 bills" so the numbers I'm using after trying different ones are 14 and 21 where there are 14 $5bills and 21 of both $10 and $20 bills

OpenStudy (anonymous):

I'ma make a small drawing and send it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ok, this is the first drawing, the little half circles represent $10 dollar bills and there are 21. ...

OpenStudy (anonymous):

At first I pretended as if there were only $10 bills, giving me a total of 210 out of the 300. Leaving me short by 90 so what I did then is I added the other half circle so that instead of the half circle= 10, the now circle=20. Now though the value is of 220, we are now off by 80 and I continued to do that until I got 300 as my final answer. Do you understand me?... I sometimes can't explain myself well, but I can try

OpenStudy (anonymous):

I understand you a little

OpenStudy (anonymous):

So not exactly right :p

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

jaja its cool :p well do you see what I did in the very beginning? how I figure out that there were 14 5 dollar bills?

OpenStudy (anonymous):

lol, yeah I get that part

OpenStudy (anonymous):

ok, well from what I meant with my drawings was this: as you can see on the first drawing: all 21 bills(half circles are worth $10) so I had said that if all 21 bills are 10s then the total value of the 35 bills that we have is 70 (from the 5's) + 210 (from the 21 10's) = 280 and leaves us short by $90 so 14 $5's 21 $10's and 0 $20's is not the bills she has. technically she has to have atleast 1 20 since she mentions she has it. so actually between the 21 bills of 10s and 20s, 1 of them is a 20 70 + 200 (20 10s) + 20( one twenty)= 290 and that leaves us short by $80... now though, since its not 370 as we like it to be there must be more 20s and less 10s so what I did I made a bill out of the 20 bills left of value $10 into 20 so now there are 19 ten dollar bills and 2 twenty dollar bills 190+40=230 plus the 70 of the 5 =300 leaving us $70 short so I did the process again. and continued to do it until I get the whole 370 i need Understand me now?

OpenStudy (anonymous):

Yes okay gotcha

OpenStudy (anonymous):

okay :D... Once you have the bills for each value you can ask me if you'd like; to see if our answers match.

OpenStudy (anonymous):

okay i get back to you

OpenStudy (anonymous):

did you get -365/2, 1227/4, -357/4 if like i did something wrong

OpenStudy (anonymous):

um..no I didn't get that.. what you do? how did you get negatives? btw, I found a shorter way to do it, based on algebra... but lets see what you did first. Could you show me your step pls

OpenStudy (anonymous):

5x+10y+20z=370, x+y+z=35, 10y+20z=7+5x some where i went wrong i guess

OpenStudy (anonymous):

um... I do think that the 10y+20z=7+5x was the problem What I was doing above was not algebra, probably elementary math lol.. and I think that if we do it that way it would look something like this: y+z+7=x

OpenStudy (anonymous):

where like you said x+y+z = 35

OpenStudy (anonymous):

but then we have many variables to solve... **** and sorry it be more like y+z-7=x

OpenStudy (anonymous):

lol it's okay

OpenStudy (anonymous):

this is what I would have done algebraically: I would have still done the whole thing I did first in trying to get how many 5s there were so there were 14 5s and a total of 21 10s and 20s so then y+z=21 and y=21-z and since 370-70=300 10y+20z=300 now, I'd plug in for the new y 10(21-z)+20z=300 then just solve for z and you'll get how many 20s there are

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

oh! and you gave me an idea of how to solve for the 5s instead of guessing let x be the dollar bills for 5's so then x+(x+7)=35 where x+7 represents the number of bills for 10s and 20s see how? one you solve for x, you still get 14

OpenStudy (anonymous):

So if you use the thing from the message right above this one, followed by the message where I solved for z, then you'll get the answer. Sorry for so many turns here and there; but finally we have arrived to the destination. Any questions?

OpenStudy (anonymous):

i get 16, 9, 10 what about you

OpenStudy (anonymous):

and thanks for your time

OpenStudy (anonymous):

uh, I got 14, 12, 9... :o..

OpenStudy (anonymous):

it can't be what you got since 9+10=19 and 19 is not 7 more than 16

OpenStudy (anonymous):

are you sloving by 3 by 3 system of equations i lead you in the wrong direction

OpenStudy (anonymous):

I don't understand what you are asking me. but basically what I solved for was for how x y and z (5s 10s and 20s)

OpenStudy (anonymous):

what i did for the last equation was set it up as x-y=7

OpenStudy (anonymous):

I see what your doing thanks lol. Sorry about that

OpenStudy (anonymous):

You kinda lost me.. lol... do you know what to do now??? and you'd be doing wrong by doing x-y=7...

OpenStudy (anonymous):

yes I shouldn't have switch and changed the signs. I see how you did it.

OpenStudy (anonymous):

I see.. well hope you got, it... I still feel confused with what you said, but I think you understood the problem with what you are writing lol...